Hewtex Electronics manufactures two products, headsets and electronic calculators, and sells them nationally to wholesalers and retailers. The Hewtex management is very pleased with the company’s performance for the current fiscal year. Projected sales suggest that 120,000 headsets and 190,000 calculators will be sold this year. The projected earnings statement which follows, shows that Hewtex will not meet its profit goal of 9% of sales after tax.

  1. Assuming that the sales mix in the plan document is achieved, how many headsets and calculator units would Hewtex have to sell to break even?

Hewtex Electronics Projected Income Statement for the year ending 31.3.2xxx

120,000 units headsets

190,000 units calculators

Total

Total Amt $000s

Per unit $

Total Amt

$ 000s

Per unit $

Amt $ 000s

Sales

1,800

15

5,320

28

7,120

Variable costs:

Materials

Labour

Others

480

360

120

4

3

1

1140

1710

570

6

9

3

1620

2070

690

Maintenance

280

2.33

1400

7.37

1680

Total costs

1240

10.33

4820

25.37

6060

Gross margin

560

4.67

500

2.63

1060

Facility sustaining cost

2000

Net income

(940)

  1. What volume of sales is required if Hewtex Electronics is to earn a profit in the next year equal to 9% of sales?
  2. Hewtex now allocates costs based on flexible labour costs. A study has determined that committed costs are as follows: (1) supervisory costs for headset production -$500,000, (2) supervisory salaries for calculator production $600,000. The balance of the committed costs are proportional to the number of batches of production. Hewtex schedules headsets for production in batch sizes of 1000, and 10,000 for calculators. Finally, $300,000 of cost which was originally seen as committed costs, was attributable to tape recorders, and $400,000 was attributable to calculators. Recast the original projection to correct the costing errors due to misclassification.
Answer #1

1.

Headsets Calculators
Selling Price per Unit $ 15 $ 28
Less: Variable Cost per Unit 8 18
Contribution Margin per Unit $ 7 $ 10
Proportion of Sales 120/310 190/310

Let the break-even quantity be Q.

At break-even point, total contribution margin equals total fixed cost.

Weighted Average Contribution Margin = 7 ( 120 / 310) Q + 10 ( 190 / 310 ) Q = 3,680

840Q + 1,900Q = 3,680 x 310 = 1,140,800

Q = 416,350 units

Number of headsets at break-even = 120 / 310 * 416,350 = 161,167.74 units

Number of calculators to break-even = 190 / 310 * 416,350 = 255,182.26 units